OpenStudy (anonymous):

what is the distance between (3,4) and (0,0)????

OpenStudy (anonymous):

Would that be the slope?

OpenStudy (anonymous):

no its distance formula

OpenStudy (amistre64):

slope isnt distance :) distance = sqrt((x-x0)^2 + (y-y0)^2) which amounts to: sqrt((3-0)^2 + (4-0)^2) in this example sqrt(3^2 + 4^2) = sqrt(9+16) = sqrt(25) = 5

OpenStudy (anonymous):

sqrt(3^2+4^2)